# Entity or Relation Embeddings? An Analysis of Encoding Strategies for Relation Extraction

### v1

@@ Abstract
Relation extraction is essentially a text classification problem, which can be tackled by fine-tuning a pre-trained language model (LM). However, a key challenge arises from the fact that relation extraction cannot straightforwardly be reduced to sequence or token classification. Existing approaches therefore solve the problem in an indirect way: they fine-tune an LM to learn embeddings of the head and tail entities, and then predict the relationship from these entity embeddings. Our hypothesis in this paper is that relation extraction models can be improved by capturing relationships in a more direct way. In particular, we experiment with appending a prompt with a [MASK] token, whose contextualised representation is treated as a relation embedding. While, on its own, this strategy significantly underperforms the aforementioned approach, we find that the resulting relation embeddings are highly complementary to what is captured by embeddings of the head and tail entity. By jointly considering both types of representations, we end up with a simple model that outperforms the state-of-the-art across several relation extraction benchmarks.
@@ 1 Introduction
Relation extraction consists in identifying the relationships between entities that are expressed in text. It is a fundamental Natural Language Processing task (NLP), which enables the learning of symbolic representations such as knowledge graphs. While Large Language Models (LLM) are highly effective in interpreting natural language inputs, the state-of-the-art in relation extraction is still based on fine-tuned language models of the BERT family Devlin et al. (2019). Moreover, relation extraction is often applied to large document collections, such as corpora of news stories, scientific articles or social media, which means that relying on LLMs is often not feasible in practice, due to their high cost.
@@
The corresponding relation embedding is obtained by concatenating the final-layer embedding of the special tokens [E1] and [E2]. This strategy was found to outperform other alternatives by Baldini Soares et al. (2019), and has remained the most popular approach for learning relation embeddings (although Zhou and Chen (2022), among others, have used type-specific special tokens, which encode the semantic type of the entities).
The success of this strategy is somewhat surprising: [E1] and [E2] are intuitively designed to represent the head and tail entities, rather than their relationship. In practice, however, knowing the semantic types of the entities often allows us to “guess” the relationship between them, especially if the types are fine-grained. For instance, knowing that the head entity is a capital city and the tail entity is a country, we may reasonably assume that the relationship being expressed is capital of. Nonetheless, we can expect that approaches based on entity embeddings may struggle to distinguish between closely related relations, especially if these relations involve entities of the same type. A possible alternative is to add a prompt which includes the [MASK] token, e.g.:
@@
We can then simply fine-tune the LM such that the contextualised embedding of the [MASK] token corresponds to a relation embedding. This strategy is popular for zero-shot and few-shot relation extraction Genest et al. (2022), but is not normally used for the standard supervised relation extraction setting. Our experiments indeed confirm that this strategy, on its own, performs poorly. Our hypothesis is that this strategy is hampered by the fact that it cannot adequately model the semantic types of the entities, which makes relation prediction much harder. The limitations of this strategy are thus complementary to those of the more common entity embedding based strategy: in one case the model almost exclusively focuses on entity types while in the other case it largely ignores them.
We present an experimental analysis which supports this hypothesis. In particular, we find that by simply combining both strategies, we can achieve substantial performance improvements, allowing us to improve the state-of-the-art in several relation extraction benchmarks. We furthermore experiment with the use of a pre-trained entity embedding model Mtumbuka and Schockaert (2023), which, despite not being trained on the relation extraction task, perform as well as the approach from Baldini Soares et al. (2019), which clearly supports our main hypothesis. As a final contribution, we also propose the use of a self-supervised pre-training strategy, which brings further performance gains.
@@ 2 Related Work
The standard approach for relation extraction with LMs uses special tokens to indicate the head and tail entities (also known as the subject and object). Such approaches compute relation embeddings by concatenating the contextualised representations of these two entities. The relation label itself is then typically predicted by training a simple classifier that takes this relation embedding as input. Several variants of this approach have been considered. The matching-the-blanks model Baldini Soares et al. (2019) encapsulates the head entity using the special tokens [E1][/E1] and the tail entity using separate special tokens [E2][/E2]. LUKE Yamada et al. (2020) simply replaces the entities by the special tokens [HEAD] and [TAIL], omitting the actual entity spans from the input. Another possibility is to use special punctuation. For instance Wang et al. (2021) encapsulate the head and tail entity with the markers @…@ and #…#, thus avoiding the introduction of new tokens. More recently, promising results have been obtained by using typed markers, which encode the semantic type of the entity. Zhong and Chen (2021) simply use type-specific special tokens, e.g. encapsulating the head entity using tokens of the form [S:type][/S:type]. However, this requires the in introduction of a large number of special tokens. Zhou and Chen (2022) avoid this by using punctuation and verbalising the semantic type as part of input. For instance, the head entity is encapsulated as @ * type *@, with type a description of the entity type.
Another possibility is to append the input with a prompt containing the [MASK] token. This strategy is popular in the context of zero-shot and few-shot relation extraction Gong and Eldardiry (2021); Chen et al. (2022); Genest et al. (2022). Rather than training a classifier on top of a relation embedding, the main aim is then to compare the contextualised representation of the MASK token with verbalisers, i.e. tokens from the language model’s vocabulary that describe the relationship. However, to the best of our knowledge, this strategy has not yet been considered in the fully supervised setting.
@@
Several approaches have been proposed for pre-training or adapting language models to make them more suitable for the task of relation extraction. The matching-the-blanks model Baldini Soares et al. (2019) uses entity linking to find sentences that refer to the same entities, and then pre-trains a relation encoder based on the idea that sentences mentioning the same entity pairs are likely to express the same relationship. More recently, several authors have studied variants of this approach based on distant supervision. Two sentences are then assumed to have the same relation if their entity pairs are asserted to have the same relation in some knowledge base, a strategy that has a long tradition in relation extraction Mintz et al. (2009). For instance, Peng et al. (2020) implement this strategy using contrastive learning with the InfoNCE loss. While the aforementioned approaches are focused on fine-tuning a pre-trained LM, the idea of changing the LM model itself has also been explored. For instance, SpanBERT Joshi et al. (2020) changes the standard masked token prediction task with the aim of learning better span-level representations, while LUKE Yamada et al. (2020) uses entity linking and distinguished entity tokens to improve the representation of entities.
@@
Most work on relation extraction uses fine-tuned models of the BERT family Devlin et al. (2019). Wan et al. (2023) discuss some of the challenges involved in using LLMs for this task, including the difficulty in selecting suitable in-context demonstrations. For this reason, the focus is still on LMs that can be fine-tuned, although recent work has started to look at models that are (considerably) larger than the standard BERT-base and BERT-large models. For instance, Sainz et al. (2021) proposed a reformulation of relation extraction as a Natural Language Inference (NLI) problem and included experiments with the 1.5B parameter DeBERTaXXL model He et al. (2021). While this allowed them to improve the state-of-the-art at the time, it should be noted that the NLI based formulation is highly inefficient when a large number of relation labels need to be considered. It also relies on manually defined verbalisations of the relation labels. Wang et al. (2022) fine-tuned a 10B parameter model on a range of tasks that can be formulated as triple prediction, including relation extraction.
@@ 3 Relation Extraction
We consider the standard sentence-level relation extraction setting, where we are given a sentence in which two entities are highlighted, which we refer to as the head entity and tail entity. The goal is to predict which relationship holds between these two entities, given a pre-defined set of candidate relation labels. We focus on strategies that first learn a relation embedding, which describes the relationship between the two entities in a continuous space, and then use a classifier to predict the actual label based on that relation embedding. In Section 3.1 we first explain the pre-training strategies that we use for learning high-quality relation embeddings. Section 3.2 then describes how the pre-trained relation encoder is fine-tuned for the relation classification task. Finally, in Section 3.3 we explain how relation embeddings can be obtained by concatenating the contextualised embeddings of the head and tail entities or by using a prompt-based strategy with a [MASK] token, among others.
@@ 3.1 Pre-training the Relation Encoder
where the temperature $\tau>0$ is a hyperparameter, and the summation in the denominator ranges over $x\in N_{s}\cup\{p\}$. The loss captures the intuitive idea that two sentences that express the same relationship should be modelled using similar relation embeddings. As suggested by previous work Baldini Soares et al. (2019); Peng et al. (2020) we also include the masked language modelling objective during pre-trainig to prevent catastrophic forgetting. The overall loss is thus given by $\mathcal{L}_{\text{info}}+\mathcal{L}_{\text{MLM}}$, with $\mathcal{L}_{\text{info}}$ the loss in (1) and $\mathcal{L}_{\text{MLM}}$ the masked language modelling objective.
@@
The effectiveness of the pre-training objective crucially depends on the quality and quantity of the available examples. In most cases, we have access to a set of labelled examples, obtained through manual annotation or distant supervision. The positive examples $P_{s}$ are then simply those examples that have the same label as $s$. As an alternative, we will also experiment with a form of self-supervised pre-training, using coreference chains as a supervision signal. Specifically, we adapt the EnCore strategy from Mtumbuka and Schockaert (2023) for pre-training entity encoders, by proposing a similar strategy for learning relation embeddings. The central idea is that two sentences are likely to express the same relationship if they refer to the same two entities. The matching-the-blank model also relies on this idea, but uses entity linking to identify such sentence pairs. Following Mtumbuka and Schockaert (2023) we select positive examples from the Gigaword corpus [^1] https://catalog.ldc.upenn.edu/LDC2003T05 and we only consider two entities to be co-referring if they are identified as such by two separate off-the-shelf coreference systems: the Explosion AI system Coreferee v1.3.1222https://github.com/explosion/coreferee and the AllenNLP coreference model [^3] https://demo.allennlp.org/coreference-resolution. This use of two coreference systems was found to reduce the number of false positives because of spurious coreference links, given that state-of-the-art coreference resolution systems are still far from perfect.
@@ 3.2 Relation Classification
Given a sentence $s$ with designated head and tail entities, we use the pre-trained relation encoder to obtain an embedding $\phi(s)$ that captures the relationship between these entities. Now consider the problem of classifying this relationship, using the labels from some set $\{l_{1},...,l_{m}\}$. Following standard practice Zhong and Chen (2021); Zhou and Chen (2022), we use a feedforward network with one hidden layer for this purpose:
@@
where $p_{i}$ is interpreted as the probability that that $l_{i}$ is the correct label, $\mathbf{A_{1}}$ and $\mathbf{A_{2}}$ are matrices, and $\mathbf{b_{1}}$ and $\mathbf{b_{2}}$ are bias terms. The label classifier is trained using cross-entropy. We also fine-tune the pre-trained relation encoder during this step.
@@ 3.3 Encoding Strategies
We now discuss the strategies we consider for obtaining relation embeddings. Suppose we are interested in the relationship between the entities <h> and <t> in sentence $s$. We first create an annotated version of sentence $s$, where (i) <h> is encapsulated with the special tokens [E1][/E1] and <t> is encapsulated with the special tokens [E2][/E2], and we append the phrase “The relation between <h> and <t> is [MASK]”. For instance, for the example from the introduction we obtain the following annotated sentence:
@@
In each of the above cases, we pre-train the relation encoder as explained in Section 3.1 and then train a classifier as explained in Section 3.2.
@@
We also consider a hybrid strategy, in which two different encoders are used during pre-training: $\phi_{1}$, which uses the [H,T] strategy, and $\phi_{2}$, which uses the Mask strategy. The pre-training loss is then of the form $\mathcal{L}_{\text{info}}^{1}+\mathcal{L}_{\text{info}}^{2}+\mathcal{L}_{\text{MLM}}$, where $\mathcal{L}_{\text{info}}^{1}$ and $\mathcal{L}_{\text{info}}^{2}$ both correspond to the InfoNCE loss (1), but where the former refers to the encoder $\phi_{1}$ and the latter refers to the encoder $\phi_{2}$. The encoders $\phi_{1}$ and $\phi_{2}$ share the same parameters, i.e. we fine-tune a single language model, but either use the concatenation of the embeddings of [E1] and [E2] as the relation embedding, or the embedding of the [MASK] token. After the pre-training step, when training the classifier, we simply concatenate the embeddings of the [E1], [E2] and [MASK] tokens. We will refer to this strategy as [H,T]+Mask. Note that the only difference with [H,T,Mask] lies in how the encoder is pre-trained. The motivation for the [H,T]+Mask strategy comes from the idea that the [MASK] token may be largely ignored when using the [H,T,Mask] strategy, since it may be easier to learn meaningful representations of the entities than to learn meaningful relation embeddings. The hybrid training strategy [H,T]+Mask avoids this issue, by forcing the [MASK] token to capture a meaningful relation embedding, before combining this representation with the contextualised entity embeddings.

### v2

@@ Abstract
Existing approaches to relation extraction obtain relation embeddings by concatenating embeddings of the head and tail entities. Despite the popularity of this approach, we find that such representations mostly capture the types of the entities involved, leading to false positives and confusion between relations that involve entities of the same type. Another possibility is to use a prompt with a [MASK] token to directly learn relation embeddings, but this approach tends to perform poorly. We show that this underperformance comes from the fact that information about entity types is insufficiently captured by the [MASK] embeddings. We therefore propose a simple model, which combines such [MASK] embeddings with entity embeddings. Despite its simplicity, our model consistently outperforms the state-of-the-art across several benchmarks, even when the entity embeddings are obtained from a pre-trained entity typing model. We also experiment with a self-supervised pre-training strategy which further improves the results. [^1] Our implementation and pre-trained models are available at https://github.com/fmtumbuka/RelationEmbeddings
@@ 1 Introduction
Relation extraction consists in identifying the relationships between entities that are expressed in text. It is a fundamental Natural Language Processing (NLP) task, which enables the learning of symbolic representations such as knowledge graphs. While large language lodels are highly effective in interpreting natural language inputs, the state-of-the-art in relation extraction is still based on fine-tuned language models of the BERT family Devlin et al. (2019). Moreover, relation extraction is often applied to large document collections, such as corpora of news stories, scientific articles or social media, which means that relying on LLMs is often not feasible in practice, due to their high cost. The study of how smaller LMs can be most effectively used for this task thus remains important.
@@
The corresponding relation embedding is obtained by concatenating the final-layer embedding of the special tokens [E1] and [E2]. This strategy was found to outperform other alternatives by Baldini Soares et al. (2019), and has remained the most popular approach for learning relation embeddings
The success of this strategy is somewhat surprising: [E1] and [E2] are intuitively designed to represent the head and tail entities, rather than their relationship even though their contextualised representations may still capture the relational context to some extent. In practice, however, knowing the semantic types of the entities often allows us to “guess” the relationship between them, especially if the types are fine-grained. For instance, knowing that the head entity is a capital city and the tail entity is a country, we may reasonably assume that the relationship being expressed is capital of. However, as our analysis in this paper shows, such approaches have at least two important limitations. First, they often struggle to distinguish relations between entities of the same type. Second, they sometimes lead to false positives (e.g. incorrectly assuming that a sentence mentioning a country and a capital city expresses the capital-of relationship). A possible alternative is to add a prompt which includes the [MASK] token, e.g.:
@@
We can then fine-tune the language model (LM) such that the contextualised embedding of [MASK] corresponds to a relation embedding. This strategy is popular for zero-shot and few-shot relation extraction Genest et al. (2022), but is not normally used for the standard supervised relation extraction setting. Our experiments indeed confirm that this approach performs poorly when used alone. Crucially, we find that this strategy struggles because it is not able to accurately characterise the semantic types of the entities, which makes relation prediction much harder. This approach is thus complementary to the entity embedding based strategy: in one case, the model nearly entirely concentrates on entity types, whereas in the other case, it mostly overlooks them. Exploiting this fact, we show that by combining both approaches, we arrive at a simple but highly effective strategy for relation extraction which improves the state-of-the-art in several relation extraction benchmarks. Our main contributions can be summarised as follows:
  • We introduce a hybrid strategy which combines the entity embedding and mask based approaches, and we empirically demonstrate its surprising effectiveness.
  • We present an analysis of the entity embedding and mask based strategies, showing that the former mostly capture the entity types while the latter do not capture the entity types to a sufficient extent. Inspired by this, we experiment with a variant in which entity embeddings from a pre-trained entity typing model Mtumbuka and Schockaert (2023) are used instead of entity embeddings that were trained for relation extraction. Surprisingly, we find that does not deteriorate the results.
  • Since the quality of entity and relation embeddings crucially depends on having access to sufficient training data, we also experiment with a self-supervised pre-training strategy and show that this strategy allows us to further improve the performance of all variants.
@@ 2 Related Work
The standard approach for relation extraction with LMs uses special tokens to indicate the head and tail entities (also known as the subject and object). Such approaches then predict relation labels from the contextualised representations of these two entities. For instance, the matching-the-blanks model Baldini Soares et al. (2019) encapsulates the head entity using the special tokens [E1][/E1] and the tail entity using separate special tokens [E2][/E2]. LUKE Yamada et al. (2020) simply replaces the entities by the special tokens [HEAD] and [TAIL], omitting the actual entity spans from the input. Wang et al. (2021) encapsulate the head and tail entity with the markers @…@ and #…#, thus avoiding the introduction of new tokens. Some approaches use typed markers, which encode the semantic type of the entity, either as special tokens Zhong and Chen (2021) or by verbalising the entity type as part of the input. It should be noted, however, that entity types are typically not available in practice, which limits the applicability of such approaches.
Another possibility is to append the input with a prompt containing the [MASK] token. This strategy is popular for zero-shot and few-shot relation extraction Gong and Eldardiry (2021); Chen et al. (2022); Genest et al. (2022). Rather than training a classifier on top of a relation embedding, the aim is then to compare the contextualised representation of the MASK token with verbalisers, i.e. tokens from the LM’s vocabulary that describe the relationship. This strategy has rarely been considered in the fully supervised setting, with KnowPrompt Chen et al. (2022) being a notable exception.
@@
Several approaches have been proposed for pre-training or adapting language models to make them more suitable for the task of relation extraction. The matching-the-blanks model Baldini Soares et al. (2019) uses entity linking to find sentences that refer to the same entities, and then pre-trains a relation encoder based on the idea that sentences mentioning the same entity pairs are likely to express the same relationship. More recently, variants of this approach based on distant supervision have also been considered. Two sentences are then assumed to have the same relation if their entity pairs are asserted to have the same relation in some knowledge base, a strategy that has a long tradition in relation extraction Mintz et al. (2009). For instance, Peng et al. (2020) implement this strategy using contrastive learning with the InfoNCE loss. While the aforementioned approaches are focused on fine-tuning a pre-trained LM, the idea of changing the LM model itself has also been explored. For instance, SpanBERT Joshi et al. (2020) changes the standard masked token prediction task with the aim of learning better span-level representations, while LUKE Yamada et al. (2020) uses entity linking and distinguished entity tokens to improve the representation of entities.
@@
LLMs such as ChatGPT perform surprisingly poorly on relation extraction benchmarks, and information extraction tasks more generally Han et al. (2023). Wan et al. (2023) discuss some of the challenges involved in using LLMs for such tasks, which include the difficulty in selecting suitable in-context demonstrations. Peng et al. (2023) make the observation that LLMs with in-context learning (ICL) struggle in particular on specification-heavy tasks, i.e. tasks where even human annotators need to carefully study a non-trivial set of annotation guidelines to correctly solve the task, as is often the case in information extraction. Due to the challenges of using ICL for relation extraction, most approaches involving LLMs use models that can be fine-tuned. For instance, Sainz et al. (2021) proposed a reformulation of relation extraction as a Natural Language Inference (NLI) problem and included experiments with the 1.5B parameter DeBERTaXXL model He et al. (2021). While this allowed them to improve the state-of-the-art at the time, it should be noted that the NLI based formulation is highly inefficient when a large number of relation labels need to be considered. It also relies on manually defined verbalisations of the relation labels. Wang et al. (2022a) fine-tuned a 10B parameter model on a range of tasks that can be formulated as triple prediction, including relation extraction. Wadhwa et al. (2023) found that while GPT-3 Brown et al. (2020) performed poorly when used directly, it was useful for generating chain-of-thought Wei et al. (2022) explanations. In particular, they showed that by fine-tuning a Flan-T5 Chung et al. (2022) model on these explanations, the resulting model performed substantially better than when fine-tuning Flan-T5 on the relation extraction task directly. As another strategy for leveraging LLMs indirectly, Xu et al. (2023) use ChatGPT for data generation in few-shot settings. Overall, however, the state-of-the-art in relation extraction, for the fully supervised setting, is still based on fine-tuned models of the BERT family Devlin et al. (2019). While this might change in future, e.g. with novel prompting techniques or better models, the need for efficient information extraction models means that such smaller models are likely to remain important.
@@ 3 Relation Extraction
We consider the standard sentence-level relation extraction setting, where we are given a sentence in which two entities are highlighted, which we refer to as the head entity and tail entity. The goal is to predict which relationship holds between these two entities, given a pre-defined set of candidate relation labels. We focus on strategies that first learn a relation embedding, which describes the relationship between the two entities in a continuous space, and then use a classifier to predict the actual label based on that relation embedding. In Section 3.1, we first explain the pre-training strategies that we use for learning high-quality relation embeddings. Section 3.2 then describes how the pre-trained relation encoder is fine-tuned for the relation classification task. Finally, in Section 3.3 we explain how relation embeddings can be obtained by concatenating the contextualised embeddings of the head and tail entities or by using a prompt-based strategy with a [MASK] token, among others.
@@ 3.1 Pre-training the Relation Encoder
where the temperature $\tau>0$ is a hyperparameter, and the summation in the denominator ranges over $x\in N_{s}\cup\{p\}$. The loss captures the intuition that two sentences expressing the same relationship should have similar relation embeddings. As suggested by previous work Baldini Soares et al. (2019); Peng et al. (2020) we also include the masked language modelling (MLM) objective during pre-training to prevent catastrophic forgetting. The overall loss is thus given by $\mathcal{L}_{\text{info}}+\mathcal{L}_{\text{MLM}}$, with $\mathcal{L}_{\text{info}}$ the loss in (1) and $\mathcal{L}_{\text{MLM}}$ the MLM objective.
@@
The effectiveness of the pre-training objective crucially depends on the quality and quantity of the available examples. In most cases, we have access to a set of labelled examples, obtained through manual annotation or distant supervision. The positive examples $P_{s}$ are then simply those examples that have the same label as $s$. As an alternative, we also experiment with a form of self-supervised pre-training, using coreference chains as a supervision signal. Specifically, we adapt the EnCore strategy from Mtumbuka and Schockaert (2023) for pre-training entity encoders, by proposing a similar strategy for learning relation embeddings. The central idea is that two sentences are likely to express the same relationship if they refer to the same two entities. The matching-the-blank model also relies on this idea, but uses entity linking to identify such sentence pairs. Following Mtumbuka and Schockaert (2023) we select positive examples from the Gigaword corpus [^2] https://catalog.ldc.upenn.edu/LDC2003T05 and we only consider two entities to be co-referring if they are identified as such by two separate off-the-shelf coreference systems: the Explosion AI system Coreferee v1.3.1333https://github.com/explosion/coreferee and the AllenNLP coreference model [^4] https://demo.allennlp.org/coreference-resolution. This use of two coreference systems was found to reduce the number of false positives because of spurious coreference links, given that state-of-the-art coreference resolution systems are still far from perfect. Clearly, the fact that two sentences mention the same entities does not guarantee that the sentences actually express the same relationship, which is a common limitation of self-supervised strategies. Note, however, that this issue is somewhat mitigated because we only consider sentences from the same news story.
@@ 3.2 Relation Classification
Given a sentence $s$ with designated head and tail entities, we use the pre-trained relation encoder to obtain an embedding $\phi(s)$ that captures the relationship between these entities. Now consider the problem of classifying this relationship, using the labels from some set $\{l_{1},...,l_{m}\}$. Following standard practice Zhong and Chen (2021); Zhou and Chen (2022), we use a feedforward network with one hidden layer and ReLU activation, i.e. predictions are made as follows:
@@
where $p_{i}$ is interpreted as the probability that that $l_{i}$ is the correct label, $\mathbf{A_{1}}$ and $\mathbf{A_{2}}$ are matrices, and $\mathbf{b_{1}}$ and $\mathbf{b_{2}}$ are bias terms. The label classifier is trained using cross-entropy. We also fine-tune the pre-trained relation encoder during this step. The dimension of the hidden representations $\mathbf{h}$ is set to be the same as that of the corresponding encoder. For instance, for all experiments with BERT-base, we set the hidden layer to 768 dimensions.
@@ 3.3 Encoding Strategies
We now discuss the considered strategies for obtaining relation embeddings. Suppose we are interested in the relationship between the entities <h> and <t> in sentence $s$. We first create an annotated version of sentence $s$, where (i) <h> is encapsulated with the special tokens [E1][/E1] and <t> is encapsulated with the special tokens [E2][/E2], and we append the phrase “The relation between <h> and <t> is [MASK]”. For instance:
@@
In each case, we pre-train the relation encoder as explained in Section 3.1 and then train a classifier as explained in Section 3.2.
Figure 1: Illustration of the [H,T]+Mask strategy.
Figure 2: Illustration of the [H,T,Mask] strategy.
@@
We also consider a hybrid approach, where we pre-train the relation encoder using a loss of the form $\mathcal{L}_{\text{info}}^{1}+\mathcal{L}_{\text{info}}^{2}+\mathcal{L}_{\text{MLM}}$. $\mathcal{L}_{\text{info}}^{1}$ and $\mathcal{L}_{\text{info}}^{2}$ both relate to the InfoNCE loss (1) but refer to different representations: $\mathcal{L}_{\text{info}}^{1}$ uses the [H,T] representation while $\mathcal{L}_{\text{info}}^{2}$ uses the Mask representation. Note that we fine-tune a single language model, i.e. the [H,T] and Mask representations are obtained with the same encoder. After the pre-training step, when training the classifier, we concatenate the two representations (i.e. the embeddings of the [E1], [E2] and [MASK] tokens). We will refer to this strategy as [H,T]+Mask. Note that the only difference with [H,T,Mask] lies in how the encoder is pre-trained. The underlying motivation comes from the idea that the [MASK] token may be largely ignored when using the [H,T,Mask] strategy, since it may be easier to learn meaningful representations of the entities than to learn meaningful relation embeddings. The hybrid training strategy [H,T]+Mask avoids this issue, by forcing the [MASK] token to capture a meaningful relation embedding, before combining this representation with the contextualised entity embeddings. Figures 1 and 2 illustrate the difference between the [H,T,Mask] and [H,T]+Mask strategies.